a) What is the probability that a “1 in a thousand” event happened sometime in the last year or 365 days?

b) What is the probability that a “1 in a thousand” event happened sometime in the last 1 thousand days?
Answer #1

probability of the event happening = 1/1000

p (not happening) = 999/1000

a) p (not happening in 365 days) = (999/1000) ^ 365 = 0.694

p(happening atleast once in 365 days) = 1 - 0.694 = 0.306

b) p (not happening in 1000 days) = (999/1000)^1000=0.367

p( happening atleast once in 1000 days) = 1 -0.367= 0.633